Can someone help me on this?
I need to solve the system by addition:
-4x - 7y = -11
x - 2y = -1
Thank you so much!
Lisa

Answer by pvaka(23)

-4x - 7y = -11
x - 2y = -1 >>Multiply by 4 >>4(x - 2y) = (-1)4 = 4x - 8y = -4
---------------------------------------------------------------
-4x - 7y = -11
+4x - 8y = -4
= -15y = -15
-15y/-15 = -15/-15
y = 1
-------------------------------------------------------------
-4x - 7(1) = -11
-4x - 7 + 7= -11 + 7
-4x = -4
-4x/-4x = -4/-4
x = 1
--------------------------------------------------------------
x = 1, y = 1